# A Simple and Effective Reinforcement Learning Method for Text-to-Image Diffusion Fine-tuning

### v4

@@ 1 Introduction
Diffusion models have emerged as a powerful tool for generative modeling (Sohl-Dickstein et al., 2015; Ho et al., 2020), with a strong capacity to model complex data distributions from various modalities, like images (Rombach et al., 2022), text (Austin et al., 2021), natural molecules (Xu et al., 2023), and videos (Blattmann et al., 2023).
@@
Diffusion models are typically pre-trained on a large-scale dataset, such that they can subsequently generate samples from the same data distribution. The training objective typically involves maximizing the data distribution likelihood. This pre-training stage helps generate high-quality samples from the model. However, some applications might require optimizing a custom reward function, for example, optimizing for generating aesthetically pleasing images (Xu et al., 2024), semantic alignment of image-text pairs based on human feedback (Schuhmann et al., 2022), or generating molecules with specific properties (Wang et al., 2024).
@@
To optimize for such black-box objectives, RL-based fine-tuning has been successfully applied to diffusion models (Fan et al., 2024; Black et al., 2023; Wallace et al., 2024; Li et al., 2024; Gu et al., 2024). For RL-based fine-tuning, the reverse diffusion process is treated as a Markov decision process (MDP), wherein prompts are treated as part of the input state, the generated image at each time-step is mapped to an action, which receives a reward from a fixed reward model (environment in standard MDP), and finally the diffusion model is treated as a policy, which we optimize to maximize rewards. For optimization, typically PPO is applied (Fan et al., 2024; Black et al., 2023). In applications where getting a reward model is infeasible or undesirable, “RL-free” fine-tuning (typically offline) can also be applied (Wallace et al., 2024; Yang et al., 2024). For this work, we only focus on diffusion model finetuning using “online” RL methods, specifically PPO (Schulman et al., 2017).
@@
Figure 1: LOOP improves attribute binding. Qualitative examples presented from images generated via Stable Diffusion (SD) 2.0 (first row), PPO (second row), and LOOP $k=4$ (third row). In the first prompt, SD and PPO both fail to bind the color black with the ball in the image, whereas LOOP binds the color black to the ball. In the second example, SD and PPO fail to generate rusted bronze color lamppost, whereas LOOP manages to do that. In the third image, SD and PPO fail to bind the shape hexagon to the watermelon, whereas LOOP manages so. In the fourth example, SD and PPO fail to generate the black horse with flowing cyan patterns, whereas LOOP generates the horse with the correct color attribute. Finally, in the last image, SD and PPO fail to bind cobalt blue color to the rock, whereas LOOP binds that successfully.
An advantage of PPO is that it removes the incentive for the new policy to deviate too much from the previous reference policy, via importance sampling and clipping operation (Schulman et al., 2017). While effective, PPO can have significant computational overhead. In practice, RL fine-tuning for diffusion models via PPO requires concurrently loading three models in memory: (i) The reference policy: The base policy, which is usually initialized with the pre-trained diffusion model. (ii) The current policy: The policy that is RL fine-tuned, and also initialized with the pre-trained diffusion model. (iii) The reward model: Typically, a large vision-language model, trained via supervised fine-tuning objective (Lee et al., 2023), which assigns a scalar reward to the final generated image during the online optimization stage. This can result in a considerable computational burden, given that each policy can potentially have millions of parameters. In addition to computational overhead, PPO is also known to be sensitive to hyper-parameters (Engstrom et al., 2019; Zheng et al., 2023; Huang et al., 2024).
Simpler approaches, like REINFORCE (Williams, 1992) avoid such complexities, and could theoretically be more efficient. However, in practice, they often suffer from high variance and instability. Recently, a variant of REINFORCE: reinforce leave-one-out (RLOO(Kool et al., 2019) was proposed which samples multiple sequences per input prompt, and a baseline correction term to reduce the variance, however, it still suffers from sample inefficiency.
This raises a fundamental question about efficiency-effectiveness trade-off in RL-based diffusion fine-tuning. In this work, first we systematically explore this trade-off between efficiency – a lower computational cost, and reduced implementation complexity (i.e., fewer hyper-parameters) – and effectiveness – stable training, and final performance. We compare a simple REINFORCE approach with the standard PPO framework, demonstrating that while REINFORCE greatly reduces computational complexity, it comes at the cost of reduced performance.
Motivated by this finding, we propose a novel RL for diffusion fine-tuning method, LOOP, which combines the best of the both worlds. To reduce the variance during policy optimization, LOOP leverages multiple actions (diffusion trajectories) and a (REINFORCE) baseline correction term per input prompt. To maintain the stability and robustness of PPO, LOOP leverages clipping and importance sampling. We note that LOOP was discovered independently in the context of long-horizon interactive LLM agent training by Chen et al. (2025).
We choose the text-to-image compositionality benchmark (T2I-CompBench; Huang et al., 2023) as our primary evaluation benchmark. Text-to-image models often fail to satisfy an essential reasoning ability of attribute binding, i.e., the generated image often fails to bind certain attributes specified in the instruction prompt (Huang et al., 2023; Ramesh et al., 2022; Fu & Cheng, 2024). As illustrated in Figure 1, LOOP outperforms previous diffusion methods on attribute binding. As attribute binding is a key skill necessary for real-world applications, we choose the T2I-CompBench benchmark alongside two other common tasks: aesthetic image generation and image-text semantic alignment.
@@
Empirical validation. To validate our claims empirically, we conduct experiments on the T2I-CompBench benchmark image compositionality benchmark. The benchmark evaluates the attribute binding capabilities of the text-to-image generative models and shows that LOOP succeeds where previous text-to-image generative models often fail. We also evaluate LOOP on two common objectives from RL for diffusion literature: image aesthetic, and text-image semantic alignment (Black et al., 2023).
@@ 2.1 Diffusion Models
We focus on denoising diffusion probabilistic models (DDPM) as the base model for text-to-image generative modeling (Ho et al., 2020; Sohl-Dickstein et al., 2015). Briefly, given a conditioning context variable $\mathbf{c}$ (text prompt in our case), and the data sample $\mathbf{x}_{0}$, DDPM models $p(\mathbf{x}_{0}\mid\mathbf{c})$ via a Markov chain of length $T$, with the following dynamics:
@@
Image generation in diffusion model is achieved via the following ancestral sampling scheme, which is a reverse diffusion process:
$\mathbf{x}_{T}{}$ $\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (2)
$\mathbf{x}_{t}$ ${}\sim N\left(\mathbf{x}_{t}|\mu_{\theta}(\mathbf{x}_{t},\mathbf{c},t),\sigma^{2}_{\theta}I\right),\forall t\in[0,T-1],$
@@ 2.2 Proximal Policy Optimization (PPO) for RL
PPO was introduced for optimizing a policy with the objective of maximizing the overall reward in the RL paradigm. PPO removes the incentive for the current policy $\pi_{t}$ to diverge from the previous policy $\pi_{t-1}$ outside the range $[1-\epsilon,1+\epsilon]$, where $\epsilon$ is a hyper-parameter. As long as the subsequent policies are closer to each other in the action space, the monotonic policy improvement bound guarantees a monotonic improvement in the policy’s performance as the optimization progresses. This property justifies the clipping term in the mathematical formulation of the PPO objective function (Schulman, 2015; Achiam et al., 2017; Queeney et al., 2021). Formally, PPO the objective function is:
$\mbox{}J(\theta)\!=\!\mathbb{E}\!\!\left[\min\!\left(\!r_{t}(\theta)\hat{A}_{t},\,\!\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\!\hat{A}_{t}\!\right)\!\right]\!,\!\mbox{}$ (3)
where $r_{t}(\theta)=\frac{\pi_{t}(a\mid c)}{\pi_{t-1}(a\mid c)}$ is the importance sampling ratio between the current policy $\pi_{t}(a\mid c)$ and the previous reference policy $\pi_{t-1}(a\mid c)$, $\hat{A}_{t}$ is the advantage function (Sutton & Barto, 2018), and the clip operator restricts the importance sampling ratio in the range $[1-\epsilon,1+\epsilon]$.
@@ 2.3 RL for Text-to-Image Diffusion Models
$\begin{split}\mathbf{s_{t}}=(\mathbf{c},t,\mathbf{x_{t}}),&\;\;\pi_{\theta}(\mathbf{a_{t}}\mid\mathbf{s_{t}})=p_{\theta}(\mathbf{x_{t-1}}\mid\mathbf{x_{t}},\mathbf{c}),\\ \mathcal{P}(\mathbf{s_{t+1}}\mid\mathbf{s_{t}},\mathbf{a_{t}})&=\delta\big{(}\mathbf{c},\mathbf{a_{t}}\big{)},\quad\mathbf{a_{t}}=\mathbf{x_{t-1}},\\ \rho_{0}(\mathbf{s_{0}})&=\big{(}p(\mathbf{c}),\delta_{T},\mathcal{N}(0,\mathbf{I})\big{)},\\ \mathcal{R}(\mathbf{s_{t}},\mathbf{a_{t}})&=\begin{cases}r(\mathbf{x_{0}},\mathbf{c})&\text{if }t=0,\\ 0&\text{otherwise.}\end{cases}\end{split}$ (4)
The input state $\mathbf{s_{t}}$ is defined in terms of the context (prompt features), sampled image at the given time-step $t$. The policy $\pi_{\theta}$ is the diffusion model itself. The state transition kernel is a dirac delta function $\delta$ with the current sampled action $\mathbf{x}_{t}$ as the input. The reward is assigned only at the last step in the reverse diffusion process, when the final image is generated. The initial state $\rho_{0}$ corresponds to the last state in the forward diffusion process: $\mathbf{x}_{T}$.
@@ 2.4 PPO for Diffusion Fine-tuning
$\begin{split}J_{\theta}(\pi)&=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[\sum_{t=0}^{T}\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\right]\\ &=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[r(\mathbf{x}_{0},\mathbf{c})\right],\end{split}$ (5)
where the trajectory $\mathbf{\tau}=\{\mathbf{x}_{T},\mathbf{x}_{T-1},\ldots,\mathbf{x}_{0}\}$ refers to the reverse diffusion process (Eq. 1), and the total reward of the trajectory is the reward of the final generated image $\mathbf{x}_{0}$ (Eq. 4). We ignore the KL-regularized version of the equation, which is commonly applied in the RLHF for LLM literature (Zhong et al., 2024; Zeng et al., 2024; Rafailov et al., 2024), and proposed by Fan et al. (2024) in the context of RL for diffusion models. As shown by Black et al. (2023), adding the KL-regularization term makes no empirical difference in terms of the final performance. The PPO objective is given as:
$J_{\theta}^{\mathrm{PPO}}(\pi)={}$ (6)
$\quad\mathbb{E}\!\bigg{[}\!\sum_{t=0}^{T}\!\text{clip}\!\left(\!\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1\!-\!\epsilon,1\!+\!\epsilon\!\right)\!r(\mathbf{x}_{0},\mathbf{c})\!\bigg{]},$
where the clipping operation removes the incentive for the new policy $\pi_{\theta}$ to differ from the previous round policy $\pi_{\text{old}}$ (Schulman et al., 2017; Black et al., 2023).
@@
REINFORCE for gradient calculation. For optimizing this objective, the REINFORCE policy gradient (also known as score function (SF)) (Williams, 1992) provides the following gradient estimate:
@@
REINFORCE with baseline correction. To reduce variance of the REINFORCE estimator, a common trick is to subtract a constant baseline correction term from the reward function (Greensmith et al., 2004; Mohamed et al., 2020):
$\begin{split}&\nabla_{\theta}J_{\theta}^{\mathrm{SFB}}(\pi)={}\\ &\quad\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}_{0},\mathbf{c}\right)-b_{t}\right)\right].\end{split}$ (9)
REINFORCE Leave-one-out (RLOO). To further reduce the variance of the REINFORCE estimator, RLOO samples $K$ diffusion trajectories per prompt ($\{\mathbf{x}^{k}_{0:T}\}\sim\pi(.\mid\mathbf{c}))$, for a better Monte-Carlo estimate of the expectation (Kool et al., 2019; Ahmadian et al., 2024). The RLOO estimator is:
$\nabla_{\theta}J_{\theta}^{\mathrm{RLOO}}(\pi)=$ (10)
$\;\mathbb{E}\left[\!K^{-1}\sum_{k=0}^{K}\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}^{k}_{t-1}\mid\mathbf{x}^{k}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}^{k}_{0},\mathbf{c}\right)-b_{t}\right)\!\right]\!.$
@@
To allow for sample reuse, the importance sampling (IS) trick can be applied (Schulman, 2015; Owen, 2013):
@@
where $\pi_{\theta}$ is the current policy to be optimized, and $\pi_{\text{old}}$ is the policy from the previous update round. With the IS trick, we can sample trajectories from the current policy in a batch, store it in a temporary buffer, and re-use them to apply mini-batch optimization (Schulman et al., 2017).
Motivation for PPO. With the IS trick, the samples from the old policy can be used to estimate the policy gradient under the current policy $\pi_{\theta}$ (Eq. 8) in a statistically unbiased fashion (Owen, 2013), i.e., in expectation the IS and REINFORCE gradients are equivalent (Eq. LABEL:eq:sf_gradient, Eq. 8). Thus, potentially, we can improve the sample efficiency of REINFORCE gradient estimation with IS.
While unbiased, the IS estimator can exhibit high variance (Owen, 2013). This high variance may lead to unstable training dynamics. Additionally, significant divergence between the current policy $\pi_{\theta}$ and the previous policy $\pi_{\textrm{old}}$ can result in the updated diffusion policy performing worse than the previous one (Schulman, 2015; Achiam et al., 2017). Next, we will prove this formally. We note that this result has previously been established by (Achiam et al., 2017) for the more general RL setting. In this work, we extend this finding to the context of diffusion model fine-tuning.
@@
$p(\mathbf{s}_{t}\mid\pi)=p(\mathbf{c})\delta(t)\!\!\int_{\mathbf{x}_{t+1}}\pi(\mathbf{x}_{t}|\mathbf{x}_{t+1},\mathbf{c},t)\pi(\mathbf{x}_{t+1}|\mathbf{c},t)\,\mathbf{dx}_{t+1}.$
@@
The advantage function is defined as: $A^{\pi_{k}}(\mathbf{s},\mathbf{a})=Q^{\pi_{k}}(\mathbf{s},\mathbf{a})-V^{\pi_{k}}(\mathbf{s})$ (Sutton & Barto, 2018). Given this, the monotonic policy improvement bound can be derived:
@@
(Achiam et al., 2017) Consider a current policy $\pi_{k}$. For any future policy $\pi$, we have:
$J(\pi)-J(\pi_{k})\geq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d^{\pi_{k}}}\left[\frac{\pi(a\mid s)}{\pi_{k}(a\mid s)}A^{\pi_{k}}(s,a)\right]$
$-\frac{2\gamma C^{\pi,\pi_{k}}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\pi_{k}}}\left[\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))\right],$
where $C^{\pi,\pi_{k}}=\max_{s\in S}|\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[A^{\pi_{k}}(s,a)\right]|$ and $\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))$ represents the total variation distance between the policies $\pi(\cdot\mid s)$ and $\pi_{k}(\cdot\mid s)$.
A direct consequence of this theorem is that when optimizing a policy with the IS objective (Eq. 11), to guarantee that the new policy will improve upon the previous policy, the policies should not diverge too much. Therefore, we need to apply a constraint on the current policy. This can be achieved by applying the clipping operator in the PPO objective (Eq. 6(Queeney et al., 2021; Achiam et al., 2017; Schulman et al., 2017; Gupta et al., 2024b).
@@
We note that a similar trade-off analysis was performed in the context of RL fine-tuning for large language models (LLM) (Ahmadian et al., 2024). However, their analysis was limited to an empirical study, whereas we present a theoretical analysis in addition to the empirical analysis. To the best of our knowledge, we are the first to conduct such a study for diffusion methods.

### v5

@@ 1 Introduction
Diffusion models have emerged as a powerful tool for generative modeling (sohl2015deep; ho2020denoising), with a strong capacity to model complex data distributions from various modalities, like images (rombach2022high), text (austin2021structured), natural molecules (xu2023geometric), and videos (blattmann2023align).
@@
Diffusion models are typically pre-trained on a large-scale dataset, such that they can subsequently generate samples from the same data distribution. The training objective typically involves maximizing the data distribution likelihood. This pre-training stage helps generate high-quality samples from the model. However, some applications might require optimizing a custom reward function, for example, optimizing for generating aesthetically pleasing images (xu2024imagereward), semantic alignment of image-text pairs based on human feedback (schuhmann2022laion), or generating molecules with specific properties (wang2024fine).
@@
To optimize for such black-box objectives, RL-based fine-tuning has been successfully applied to diffusion models (fan2024reinforcement; black2023training; wallace2024diffusion; li2024aligning; gu2024diffusion). For RL-based fine-tuning, the reverse diffusion process is treated as a Markov decision process (MDP), wherein prompts are treated as part of the input state, the generated image at each time-step is mapped to an action, which receives a reward from a fixed reward model (environment in standard MDP), and finally the diffusion model is treated as a policy, which we optimize to maximize rewards. For optimization, typically PPO is applied (fan2024reinforcement; black2023training). In applications where getting a reward model is infeasible or undesirable, “RL-free” fine-tuning (typically offline) can also be applied (wallace2024diffusion). For this work, we only focus on diffusion model fine-tuning using “online” RL methods, specifically PPO (schulman2017proximal).
An advantage of PPO is that it removes the incentive for the new policy to deviate too much from the previous reference policy, via importance sampling and clipping operation (schulman2017proximal). While effective, PPO can have significant computational overhead. In practice, RL fine-tuning for diffusion models via PPO requires concurrently loading three models in memory: (i) The reference policy: The base policy, which is usually initialized with the pre-trained diffusion model. (ii) The current policy: The policy that is RL fine-tuned, and also initialized with the pre-trained diffusion model. (iii) The reward model: Typically, a large vision-language model, trained via supervised fine-tuning objective (lee2023aligning), which assigns a scalar reward to the final generated image during the online optimization stage. This can result in a considerable computational burden, given that each policy can potentially have millions of parameters. In addition to computational overhead, PPO is also known to be sensitive to hyper-parameters (engstrom2019implementation; zheng2023delve; huang2024n+).
Simpler approaches, like REINFORCE (williams1992simple) avoid such complexities, and could theoretically be more efficient. However, in practice, they often suffer from high variance and instability. A variant of REINFORCE: reinforce leave-one-out (RLOO(kool2019buy) has been proposed that samples multiple sequences per input prompt, and a baseline correction term to reduce the variance; however, it still suffers from sample inefficiency.
This raises a fundamental question about the efficiency-effectiveness trade-off in RL-based diffusion fine-tuning. In this work, first we systematically explore this trade-off between efficiency – a lower computational cost, and reduced implementation complexity (i.e., fewer hyper-parameters) – and effectiveness – stable training, and final performance. We compare a simple REINFORCE approach with the standard PPO framework, demonstrating that while REINFORCE greatly reduces computational complexity, it comes at the cost of reduced performance.
Motivated by this finding, we propose a novel RL for diffusion fine-tuning method, LOOP, which combines the best of the both worlds. To reduce the variance during policy optimization, LOOP leverages multiple actions (diffusion trajectories) and a (REINFORCE) baseline correction term per input prompt. To maintain the stability and robustness of PPO, LOOP leverages clipping and importance sampling.
Our approach is conceptually similar to the recently proposed GRPO method for RL fine-tuning of LLMs (shao2024deepseekmath). The key technical differences are: (i) LOOP does not apply standard-deviation normalization in the advantage calculation. Recent work on LLM fine-tuning suggests that removing this normalization term can improve performance (liu2025understanding). (ii) Following this recent work, LOOP omits the KL penalty term. Prior studies indicate that explicit KL regularization has minimal practical effect on performance (Black2023), and recent theoretical work shows that on-policy RL methods implicitly maintain KL proximity to the base policy even without explicit regularization (shenfeld2025rl). (iii) In the diffusion setting, the reverse process has a fixed sequence length across all generations, making sequence-length normalization unnecessary.
For the primary evaluation benchmark, we choose the text-to-image compositionality benchmark (T2I-CompBench; huang2023t2i. Text-to-image models often fail to satisfy an essential reasoning ability of attribute binding, i.e., the generated image often fails to bind certain attributes specified in the instruction prompt (huang2023t2i; ramesh2022hierarchical; fu2024enhancing). As illustrated in Figure 1, LOOP outperforms previous diffusion methods on attribute binding. As attribute binding is a key skill necessary for real-world applications, we choose the T2I-CompBench benchmark alongside two other common tasks: aesthetic image generation and image-text semantic alignment.
@@
DDPO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 1: LOOP improves attribute binding. Qualitative examples presented from images generated via Stable Diffusion (SD) 2.0 (first row), DDPO (black2023training) (second row), and LOOP $k=4$ (third row). In the first prompt, SD and DDPO both fail to bind the color black with the ball in the image, whereas LOOP binds the color black to the ball. In the second example, SD and DDPO fail to generate rusted bronze color lamppost, whereas LOOP manages to do that. In the third image, SD and DDPO fail to bind the shape hexagon to the watermelon, whereas LOOP manages so. In the fourth example, SD and DDPO fail to generate the black horse with flowing cyan patterns, whereas LOOP generates the horse with the correct color attribute. Finally, in the last image, SD and DDPO fail to bind cobalt blue color to the rock, whereas LOOP binds that successfully.
@@
  • Empirical validation. To validate our claims empirically, we conduct experiments on the T2I-CompBench benchmark image compositionality benchmark. The benchmark evaluates the attribute binding capabilities of the text-to-image generative models and shows that LOOP succeeds where previous text-to-image generative models often fail. We also evaluate LOOP on two common objectives from the literature on RL for diffusion: image aesthetic and text-image semantic alignment (black2023training).
The remainder of the paper is organized as follows. In the next section, we provide the necessary background and discuss related work. Section 3 revisits the efficiency–effectiveness trade-off between REINFORCE and PPO. Section 4 introduces our proposed method, Leave-One-Out PPO (LOOP) for diffusion fine-tuning. Section 5 describes the experimental setup, and Section 6 presents the results and discussion. Finally, Section 7 concludes the paper.
@@ 2.1 Diffusion Models
We focus on denoising diffusion probabilistic models (DDPM) as the base model for text-to-image generative modeling (ho2020denoising; sohl2015deep). Briefly, given a conditioning context variable $\mathbf{c}$ (a text prompt in our case), and the data sample $\mathbf{x}_{0}$, DDPM models $p(\mathbf{x}_{0}\mid\mathbf{c})$ via a Markov chain of length $T$, with the following dynamics:
@@
Image generation in a diffusion model is achieved via the following ancestral sampling scheme, which is a reverse diffusion process:
$\mathbf{x}_{T}{}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\;\;\mathbf{x}_{t}\sim N\left(\mathbf{x}_{t}|\mu_{\theta}(\mathbf{x}_{t},\mathbf{c},t),\sigma^{2}_{\theta}I\right),\forall t\in[0,T-1],$ (2)
@@ 2.2 Proximal Policy Optimization (PPO) for RL
PPO was introduced for optimizing a policy with the objective of maximizing the overall reward in the RL paradigm. PPO removes the incentive for the current policy $\pi_{t}$ to diverge from the previous policy $\pi_{t-1}$ outside the range $1-\epsilon,1+\epsilon]$, where $\epsilon$ is a hyper-parameter. As long as the subsequent policies are closer to each other in the action space, the monotonic policy improvement bound guarantees a monotonic improvement in the policy’s performance as the optimization progresses. This property justifies the clipping term in the mathematical formulation of the PPO objective function (schulman2015trust; achiam2017constrained; queeney2021generalized). Formally, the PPO objective function is:
$\mbox{}\hskip-5.69054ptJ(\theta)\!=\!\mathbb{E}\!\!\left[\min\!\left(\!r_{t}(\theta)\hat{A}_{t},\,\!\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\!\hat{A}_{t}\!\right)\!\right]\!,\!\mbox{}$ (3)
where $r_{t}(\theta)=\frac{\pi_{t}(a\mid c)}{\pi_{t-1}(a\mid c)}$ is the importance sampling ratio between the current policy $\pi_{t}(a\mid c)$ and the previous reference policy $\pi_{t-1}(a\mid c)$, $\hat{A}_{t}$ is the advantage function (sutton2018reinforcement), and the clip operator restricts the importance sampling ratio in the range $1-\epsilon,1+\epsilon]$.
@@ 2.3 RL for Text-to-Image Diffusion Models
$\begin{split}\mathbf{s_{t}}=(\mathbf{c},t,\mathbf{x_{t}}),\;\;\pi_{\theta}(\mathbf{a_{t}}\mid\mathbf{s_{t}})=p_{\theta}(\mathbf{x_{t-1}}\mid\mathbf{x_{t}},\mathbf{c}),\;\;\mathcal{P}(\mathbf{s_{t+1}}\mid\mathbf{s_{t}},\mathbf{a_{t}})= \delta\big(\mathbf{c},\mathbf{a_{t}}\big),\;\;\mathbf{a_{t}}=\mathbf{x_{t-1}},\\ \rho_{0}(\mathbf{s_{0}})=\big(p(\mathbf{c}),\delta_{T},\mathcal{N}(0,\mathbf{I})\big),\;\;\mathcal{R}(\mathbf{s_{t}},\mathbf{a_{t}})= \begin{cases}r(\mathbf{x_{0}},\mathbf{c})&\text{if }t=0,\\ 0&\text{otherwise.}\end{cases}\end{split}$ (4)
The input state $\mathbf{s_{t}}$ is defined in terms of the context $\mathbf{c}$ (prompt features), and the sampled image at the given time-step $t$: $\mathbf{x_{t}}$. The policy $\pi_{\theta}$ is the diffusion model itself. The state transition kernel is a dirac delta function $\delta$ with the current sampled action $\mathbf{x}_{t}$ as the input. The reward is assigned only at the last step in the reverse diffusion process, when the final image is generated. The initial state $\rho_{0}$ corresponds to the last state in the forward diffusion process: $\mathbf{x}_{T}$.
@@ 2.4 PPO for Diffusion Fine-tuning
$J_{\theta}(\pi)=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[\sum_{t=0}^{T}\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\right]=\mathbb{E}_{\tau\sim p(\tau\mid\pi_{\theta})}\left[r(\mathbf{x}_{0},\mathbf{c})\right],$ (5)
where the trajectory $\mathbf{\tau}=\{\mathbf{x}_{T},\mathbf{x}_{T-1},\ldots,\mathbf{x}_{0}\}$ refers to the reverse diffusion process (Eq. 1), and the total reward of the trajectory is the reward of the final generated image $\mathbf{x}_{0}$ (Eq. 4). We ignore the KL-regularized version of the equation, which is commonly applied in the RLHF for LLM literature (zhong2024dpo; zeng2024token; rafailov2024direct), and proposed by fan2024reinforcement in the context of RL for diffusion models. As shown by black2023training, adding the KL-regularization term makes no empirical difference in terms of the final performance. The PPO objective is given as:
$J_{\theta}^{\mathrm{PPO}}(\pi)=\mathbb{E}\!\bigg[\!\sum_{t=0}^{T}\!\text{clip}\!\left(\!\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1\!-\!\epsilon,1\!+\!\epsilon\!\right)\!r(\mathbf{x}_{0},\mathbf{c})\!\bigg],$
where the clipping operation removes the incentive for the new policy $\pi_{\theta}$ to differ from the previous round policy $\pi_{\text{old}}$ (schulman2017proximal; black2023training).
@@
REINFORCE for gradient calculation. For optimizing this objective, the REINFORCE policy gradient (also known as score function (SF)) (williams1992simple) provides the following gradient estimate:
@@
REINFORCE with baseline correction. To reduce variance of the REINFORCE estimator, a common trick is to subtract a constant baseline correction term from the reward function (greensmith2004variance; mohamed2020monte):
$\nabla_{\theta}J_{\theta}^{\mathrm{SFB}}(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}_{0},\mathbf{c}\right)-b_{t}\right)\right].$ (8)
REINFORCE Leave-one-out (RLOO). To further reduce the variance of the REINFORCE estimator, RLOO samples $K$ diffusion trajectories per prompt ($\{\mathbf{x}^{k}_{0:T}\}\sim\pi(.\mid\mathbf{c}))$, for a better Monte-Carlo estimate of the expectation (kool2019buy; ahmadian2024back). The RLOO estimator is:
$\nabla_{\theta}J_{\theta}^{\mathrm{RLOO}}(\pi)=\mathbb{E}\left[\!K^{-1}\sum_{k=0}^{K}\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}\left(\mathbf{x}^{k}_{t-1}\mid\mathbf{x}^{k}_{t},\mathbf{c}\right)\left(r\left(\mathbf{x}^{k}_{0},\mathbf{c}\right)-b_{t}\right)\!\right]\!.\\$ (9)
@@
To allow for sample reuse, the importance sampling (IS) trick can be applied (schulman2015trust; mcbook):
@@
where $\pi_{\theta}$ is the current policy to be optimized, and $\pi_{\text{old}}$ is the policy from the previous update round. With the IS trick, we can sample trajectories from the current policy in a batch, store it in a temporary buffer, and re-use them to apply mini-batch optimization (schulman2017proximal).
Motivation for PPO. With the IS trick, the samples from the old policy can be used to estimate the policy gradient under the current policy $\pi_{\theta}$ (Eq. 7) in a statistically unbiased fashion (mcbook), i.e., in expectation the IS and REINFORCE gradients are equivalent (Eq. 10, Eq. 7). Thus, potentially, we can improve the sample efficiency of REINFORCE gradient estimation with IS.
While unbiased, the IS estimator can exhibit high variance (mcbook). This high variance may lead to unstable training dynamics. Additionally, significant divergence between the current policy $\pi_{\theta}$ and the previous policy $\pi_{\textrm{old}}$ can result in the updated diffusion policy performing worse than the previous one (schulman2015trust; achiam2017constrained). Next, we will prove this formally. We note that this result has previously been established by (achiam2017constrained) for the more general RL setting. In this work, we extend this finding to the context of diffusion model fine-tuning.
@@
$p(\mathbf{s}_{t}\mid\pi)=p(\mathbf{c})\delta(t)\!\!\int_{\mathbf{x}_{t+1}}\hskip-15.6491pt\pi(\mathbf{x}_{t}|\mathbf{x}_{t+1},\mathbf{c},t)\pi(\mathbf{x}_{t+1}|\mathbf{c},t)\,\mathbf{dx}_{t+1}.$
@@
The advantage function is defined as: $A^{\pi_{k}}(\mathbf{s},\mathbf{a})=Q^{\pi_{k}}(\mathbf{s},\mathbf{a})-V^{\pi_{k}}(\mathbf{s})$ (sutton2018reinforcement). Given this, the monotonic policy improvement bound can be derived:
@@
(achiam2017constrained) Consider a current policy $\pi_{k}$. Let $C^{\pi,\pi_{k}}=\max_{s\in S}|\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[A^{\pi_{k}}(s,a)\right]|$, and $\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))$ represent the total variation distance between the policies $\pi(\cdot\mid s)$ and $\pi_{k}(\cdot\mid s)$, and $s$ be the current state. For any future policy $\pi$, we have:
$J(\pi)-J(\pi_{k})\geq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d^{\pi_{k}}}\left[\frac{\pi(a\mid s)}{\pi_{k}(a\mid s)}A^{\pi_{k}}(s,a)\right]-\frac{2\gamma C^{\pi,\pi_{k}}}{(1-\gamma)^{2}}\mathbb{E}_{s\sim d^{\pi_{k}}}\left[\mathrm{TV}(\pi(\cdot\mid s),\pi_{k}(\cdot\mid s))\right].$
A direct consequence of this theorem is that when optimizing a policy with the IS objective (Eq. 10), to guarantee that the new policy will improve upon the previous policy, the policies should not diverge too much. Therefore, we need to apply a constraint on the current policy. This can be achieved by applying the clipping operator in the PPO objective (Eq. 3(queeney2021generalized; achiam2017constrained; schulman2017proximal; gupta2024practical; gupta2024unbiased).
@@
We note that a similar trade-off analysis was performed in the context of RL fine-tuning for large language models (LLM) (ahmadian2024back). However, their analysis was limited to an empirical study, whereas we present a theoretical analysis in addition to the empirical analysis. To the best of our knowledge, we are the first to conduct such a study for diffusion methods.